# Rethinking 3D Dense Caption and Visual Grounding in A Unified Framework through Prompt-based Localization

### v1

@@ 3.4 Caption Text Prompt
BUTD-DETR (Jain et al., 2022) suggests that the task of object detection is essentially a form of referential language grounding, where the utterance is simply the object’s category label. Essentially, object detection can be viewed as the grounding process of detection prompts.
Concretely, utilizing the detector’s library of object categories, we form prompts by stringing together the labels of objects to be identified, like “cabinet . bed . chair . sofa . ”. These concatenated labels are treated as reference text for grounding: the goal is to locate all instances of the mentioned categories in the scene, if present. Additionally, incorporating negative category labels (for which no instances exist in the scene) serves as negative training, teaching the model to avoid matching any boxes to these negative labels.
@@ 3.5 Multi-task End-to-end Training
For the 3DVG task, we follow the training scheme of EDA (Wu et al., 2023), which adopts five loss functions for each layer of the object decoder: box center coordinate prediction with a smooth-$L_{1}$ loss $\mathcal{L}_{\text{coord }}$, box size prediction with a smooth-$L_{1}$ loss $\mathcal{L}_{\text{size }}$, GIoU loss (Rezatofighi et al., 2019) $\mathcal{L}_{\text{giou }}$, the semantic alignment loss $\mathcal{L}_{\text{sem }}$, and the position alignment loss $\mathcal{L}_{\text{pos }}$. The loss of $l$-th decoder layer is the combination of these 5 loss terms by weighted summation:
@@
The losses on all decoder layers are averaged to form the total 3DVG loss:
$\mathcal{L}_{\text{vg }}=\frac{1}{L}\sum_{l=1}^{L}\mathcal{L}_{\text{dec }}^{(l)}.$
For the 3DDC task, we follow Vote2Cap-DETR (Chen et al., 2023a) and apply the standard cross-entropy loss (MLE training), then fine-tune it with Self-Critical Sequence Training (SCST). We adopt the standard SCST in our model, whose reward function is CIDEr (Vedantam et al., 2015) score. We termed the caption loss as $\mathcal{L}_{\text{cap }}$. During MLE training, together with the KPS loss (Liu et al., 2021) $\mathcal{L}_{\text{kps }}$ for the query selection, the final loss for 3DGCTR in end-to-end manner is as follows:
@@
Note that when the batch of training data is 3DVG, $\mathcal{L}_{\text{cap }}=0$. While in SCST training, the VG model is frozen and only the caption head is trained with the caption loss $\mathcal{L}_{\text{cap }}$:
$\mathcal{L}=\mathcal{L}_{\text{cap}}.$
We also use the set-to-set training strategy proposed by Vote2Cap-DERT (Chen et al., 2023a) for better performance.
@@ 3.6 Inference
During inference, given an input point cloud and a referring text, 3DGCTR outputs the object boxes $B\in\mathbb{R}^{k\times 6}$. Subsequently, the referent box will be selected according to the referring scores $s\in\mathbb{R}^{k}$. Specifically, in 3DVG, the target box described in the referring text will be selected, while in 3DDC, all label boxes described in the prompt are selected, and the caption head provides a dense caption for each box.
@@
ScanRefer (Chen and Chang, 2020) 67.64 49.19 32.06 21.26 38.97 26.10
TGNN (Huang et al., 2021) 68.61 56.80 29.84 23.18 37.37 29.70
3DJCG (Cai et al., 2022) 78.75 61.30 40.13 30.08 47.62 36.14
D3Net (Chen et al., 2021a) - 70.35 - 30.50 - 37.87
UniT3D (Chen et al., 2023b) 82.75 73.14 36.36 31.05 45.27 39.14
3DSPS (Luo et al., 2022) 84.12 66.72 40.32 29.82 48.82 36.98
BUTD-DETR (Jain et al., 2022) 82.88 64.98 44.73 33.97 50.42 38.60
EDA (Wu et al., 2023) 85.76 68.57 49.13 37.64 54.59 42.26
EDA-PMB (ours backbone) 88.22 73.40 51.76 40.82 57.45 45.91
@@
Table 1: The 3D referring expression comprehension results on ScanRefer in terms of Acc@0.25IoU and Acc@0.5IoU. Note that what is shown in the table is the optimal accuracy of each method as reported in its paper. Our 3DGCTR surpasses existing state-of-the-art works by a significant margin. EDA-PMB refers to the replacement of single-stage EDA’s (Wu et al., 2023) visual backbone from Pointnet++ (Qi et al., 2017) to PointMetaBase (Lin et al., 2023).
4 Experiments
4.1 DataSets and Metrics
We assess our approach on 3D referring datasets ScanRefer (Chen and Chang, 2020). ScanRefer derives from ScanNet (Dai et al., 2017), featuring 1,513 detailed 3D in-door scene reconstructions. Specifically, ScanNet (Dai et al., 2017) comprises 1,201 training and 312 validation indoor 3D scenes. ScanRefer adheres to the official ScanNet splits, it offers 51,583 descriptions, averaging 13.81 objects and 64.48 descriptions per scene.
In terms of the evaluation metric, for 3DVG, we use Acc@IoU to measure the proportion of descriptions where the predicted box and ground truth overlap with an IoU greater than 0.25 and 0.5. Descriptions are categorized into “unique” if the object is the sole representative of its class in the scene, or “multiple” otherwise. As for 3DDC, following (Chen et al., 2023a), the evaluation metric is $m$@$k$IoU that $m$ could be any metric for natural language generation, such as CIDEr (Vedantam et al., 2015), METEOR (Banerjee and Lavie, 2005), BLEU-4 (Papineni et al., 2002), and ROUGE-L (Lin, 2004). In practice, $k$ is set to 0.25 and 0.5.
4.2 Implementation Details
We first pre-train the 3DVG model without the caption head on ScanRefer datasets (Chen and Chang, 2020) following the training settings in (Wu et al., 2023). During MLE training, we load the pre-trained 3DVG weight and jointly train the whole network for another 30 epochs in both VG and DC tasks. To prevent overfitting, the initial learning rates of the 3DVG model and caption head are empirically set to 2e-6 and 2e-4, respectively. We apply learning rate decay at epochs 10 and 20 with a rate of 0.1. As for the input type, we use XYZ coordinates and RGB values as the input, and the number of visual tokens $V$ and query embeddings $Q$ are empirically set to 1024 and 256, respectively. The number of decoder layers $L$ is set to 6. The balancing factors are set default as $\alpha_{1}$ = 1.0 / ($L$ + 1), $\alpha_{2}$ = 5.0, $\alpha_{3}$ = 8.0, $\beta_{1}$ = 5.0, $\beta_{2}$ = 1.0, $\beta_{3}$ = 1.0, $\beta_{4}$ = 0.5 and $\beta_{5}$ = 0.5 for the ScanRefer dataset. During SCST training, due to the high memory cost, we tune the caption head with a batch size of 8 and freeze the rest of the modules for 400 epochs with a fixed learning rate of 5e-6 and apply learning rate decay at epochs 100 and 200 with a rate of 0.1.
@@ 4.2 Implementation Details
Scan3Cap (Chen et al., 2021b) 56.82 34.18 26.29 55.27 39.08 23.32 21.97 44.78
MORE (Jiao et al., 2022) 62.91 36.25 26.75 56.33 40.94 22.93 21.66 44.42
SpaCap3d (Wang et al., 2022a) 63.30 36.46 26.71 55.71 44.02 25.26 22.33 45.36
3DJCG (Cai et al., 2022) 64.70 40.17 27.66 59.23 49.48 31.03 24.22 50.80
D3Net (Chen et al., 2021a) - - - - 46.07 30.29 24.35 51.67
UniT3D (Chen et al., 2023b) - - - - 46.69 27.22 21.91 45.98
Vote2Cap-DETR (Chen et al., 2023a) 71.45 39.34 28.25 59.33 61.81 34.46 26.22 54.40
@@
$\mathcal{X}$-Trans2Cap (Yuan et al., 2022) 61.83 35.65 26.61 54.70 43.87 25.05 22.46 45.28
Scan3Cap (Chen et al., 2021b) - - - - 48.38 26.09 22.15 44.74
D3Net (Chen et al., 2021a) - - - - 62.64 35.68 25.72 53.90
Vote2Cap-DETR (Chen et al., 2023a) 84.15 42.51 28.47 59.26 73.77 38.21 26.64 54.71
@@
4.3 Quantitative Comparison
In this subsection, we perform a qualitative comparison on ScanRefer (Chen and Chang, 2020) datasets.
3DVG. To evaluate the 3DVG performance, we compare 3DGCTR with several existing 3DVG works in the 3D Referring Expression Comprehension (3DREC) task on ScanRefer, involving the state-of-the-art DETR-like methods EDA (Wu et al., 2023) and BUTD-DETR (Jain et al., 2022), among others. As evidenced in Table. 1, 3DGCTR consistently exhibits superior performance over these prevailing methods in both two-stage and single-stage settings. This evidence strongly suggests that the newly incorporated lightweight caption head has indeed augmented the performance of the 3DVG task.
3DDC. In 3DDC performance, among the evaluated methods, most employ the standard VoteNet (Ding et al., 2019) detector, except D3Net (Chen et al., 2021a) and 3DJCG (Cai et al., 2022). The state-of-the-art method Vote2Cap-DETR (Chen et al., 2023a) is the first attempt to jointly train 3D Detection and 3D Dense Caption tasks in a single-stage manner, which gains state-of-the-art performance. Table. 2 reports comparisons on ScanRefer (Chen and Chang, 2020) dataset. Our 3DGCTR surpasses current state-of-the-art methods. For example, our 3DGCTR achieves 66.11% C@0.5 while Vote2Cap-DETR. (Chen et al., 2023a) achieves 61.81% (4.30% C@0.5↑). Additionally, under SCST, our 3DGCTR outperforms the state-of-the-art method C@0.5 (0.18% C@0.5↑).
Figure 3: Qualitative Comparisons. We compare qualitative results with two state-of-the-art methods in 3DVG (left part in the figure) and 3DDC (right part in the figure) tasks, EDA (Wu et al., 2023) and Vote2Cap-DETR (Chen et al., 2023a). We mark correct attribute words in green and wrong descriptions in red. Our method produces right bounding boxes close to ground truth annotations and produces accurate descriptions of object attributes, classes and spatial relationships.
@@ 4.4 Ablation Study
Influence of 3DVG model. We replace the pre-trained 3DVG model with EDA (Wu et al., 2023) in both single- and two-stage settings. As illustrated in Table. 3, under the joint training of 3DDC and 3DVG, various 3DVG backbones achieve enhanced performance in 3DVG tasks (Acc@0.25IoU). Meanwhile, when using the state-of-the-art 3DVG model (Wu et al., 2023) as the component of our method, we obtain the state-of-the-art 3DDC performance compared to Vote2Cap-DETR (Chen et al., 2023a) (3.29% C@0.5↑), which demonstrate that our “3DDC builds upon 3DVG model” rethinking pipeline offers greater advantages.
@@
Vote2Cap (Chen et al., 2023a) 61.81 -
EDAs (Wu et al., 2023) - 53.83
@@
Table 3: Performance on Scanrefer affected by different 3DVG backbone during MLE training. The expressions EDAs and EDAt denote the application of EDA (Wu et al., 2023) in single- and two-stage settings, respectively. The acronyms 3DGCTR-EDAs and 3DGCTR-EDAt are used to indicate the utilization of EDAs and EDAt as the 3DVG backbone in the corresponding settings. The red font represents the performance improvement after joint training.
Joint Training Scheme. We explore the impacts of different joint training schemes on task performance. After pre-training the 3DVG model without the caption head, if only dense caption data is used for training, the training scheme can be divided into the following three types: 1) only 3DDC: training all components with the same learning rate, 2) only 3DDC (two lr): training all components with different learning rates ($42e-$ 4 and $62e-$ 6), 3) only 3DDC (frozen vg): freezing 3DVG components and training 3DDC components. If both data to jointly train 3DDC and 3DVG tasks, the training scheme can be divided into the following two types: 4) joint training: jointly training 3DDC and 3DVG with the same learning rate, 5) joint training (two lr): jointly training 3DDC and 3DVG with different learning rates ($42e-$ 4 and $62e-$ 6), As displayed in Table. 4, The first three schemes bring about significant 3DDC performance gain but cause a suboptimal 3DVG performance. Using the same learning rate in the first scheme significantly decreases 3DVG performance, as solely employing 3DDC training data biases the model’s statistical parameters towards DC data (e.g., mean as well as variation in the batch normalization layer), negatively impacting 3DVG. This is mitigated in the second and third schemes by either freezing or reducing the learning rate for 3DVG components, which preserves 3DVG localization and improves 3DDC performance. However, applying the same learning rate for both 3DDC and 3DVG data in the fourth scheme leads to overfitting in 3DVG training, skewing the model’s statistical parameters towards VG data and impairing 3DDC performance. In contrast, the fifth scheme, which uses different learning rates, achieves optimal 3DDC accuracy and slightly better 3DVG performance. It can also be seen from the table that after joint training, 3DGCTR’s 3DVG performance increased by 0.3% compared with EDA-PMB, and 3DDC performance increased by 1.27% compared with only 3DDC training, suggesting that a well-designed training approach allows for effective end-to-end training of both 3DDC and 3DVG tasks within the same framework, benefiting each other.
@@
5) joint training (two lr) 66.11(+1.27%) 57.75(+0.30%)
Table 4: Performance on Scanrefer affected by different training schemes during MLE training. ’two lr’ means jointly training 3DDC and 3DVG components with different learning rates, and ’frozen vg’ means only training 3DDC components while freezing the 3DVG backbone. The red font represents the performance improvement after joint training, while underline represents the performance to be compared with our optimal precision.

### v2

@@ 3.4 Caption Text Prompt
BUTD-DETR [17] suggests that the task of object detection is essentially a form of referential language grounding, where the utterance is simply the object’s category label. Essentially, object detection can be viewed as the grounding process of detection prompts.
Concretely, utilizing the detector’s library of object categories, we form prompts by stringing together the labels of objects to be identified, like “cabinet . bed . chair . sofa . ". These concatenated labels are treated as reference text for grounding: the goal is to locate all instances of the mentioned categories in the scene, if present. Additionally, incorporating negative category labels (for which no instances exist in the scene) serves as negative training, teaching the model to avoid matching any boxes to these negative labels.
@@ 3.5 Multi-task End-to-end Training
For the 3DVG task, we follow the training scheme of EDA [34], which adopts five loss functions for each layer of the object decoder: box center coordinate prediction with a smooth-$L_{1}$ loss $\mathcal{L}_{\text{coord }}$, box size prediction with a smooth-$L_{1}$ loss $\mathcal{L}_{\text{size }}$, GIoU loss [28] $\mathcal{L}_{\text{giou }}$, the semantic alignment loss $\mathcal{L}_{\text{sem }}$, and the position alignment loss $\mathcal{L}_{\text{pos }}$. The loss of $l$-th decoder layer is the combination of these 5 loss terms by weighted summation:
@@
The losses on all decoder layers are averaged to form the total 3DVG loss: $\mathcal{L}_{\text{vg }}=\frac{1}{L}\sum_{l=1}^{L}\mathcal{L}_{\text{dec }}^{(l)}$.
For the 3DDC task, we follow Vote2Cap-DETR [8] and apply the standard cross-entropy loss (MLE training), then fine-tune it with Self-Critical Sequence Training (SCST). We adopt the standard SCST in our model, whose reward function is CIDEr [31] score. We termed the caption loss as $\mathcal{L}_{\text{cap }}$. During MLE training, together with the KPS loss [23] $\mathcal{L}_{\text{kps }}$ for the query selection, the final loss for 3DGCTR in end-to-end manner is as follows:
@@ 3.6 Inference
During inference, given an input point cloud and a referring text, 3DGCTR outputs the object boxes $B\in\mathbb{R}^{k\times 6}$. Subsequently, the referent box will be selected according to the referring scores $s\in\mathbb{R}^{k}$. In 3DVG, the target box described in the referring text will be selected, while in 3DDC, all boxes described in the prompt are selected, and the caption head provides a dense caption for each box.
4 Experiments
4.1 DataSets and Metrics
We utilized two widely recognized datasets: ScanRefer [5] and Nr3D [1], both of which are based on ScanNet’s [11] 3D indoor scenes. Specifically, ScanNet [11] comprises 1,201 training and 312 validation indoor 3D scenes. The ScanRefer [5] and Nr3D [1] datasets include 36,665 and 32,919 free-form language descriptions, respectively, referencing 7,875 and 4,664 objects within 562 and 511 training scenes. The evaluation was conducted on 9,508 and 8,584 sentences corresponding to 2,068 and 1,214 objects across 141 and 130 3D scenes, respectively.
In terms of the evaluation metric, for 3DVG, we use Acc@IoU to measure the proportion of descriptions where the predicted box and ground truth overlap with an IoU greater than 0.25 and 0.5. Descriptions are categorized into “unique" if the object is the sole representative of its class in the scene, or “multiple" otherwise. As for 3DDC, following [8], the evaluation metric is $m$@$k$IoU that $m$ could be any metric for natural language generation, such as CIDEr [31], METEOR [2], BLEU-4 [25], and ROUGE-L [20]. In practice, $k$ is set to 0.25 and 0.5.
4.2 Implementation Details
We first pre-train the 3DVG model without the caption head on ScanRefer [5] and Nr3D[1] datasets following the training settings in [34] on four NVIDIA V100 GPUs for 4 days. During MLE training, we load the pre-trained 3DVG weight and jointly train the whole network for another 30 epochs in both VG and DC tasks for 12 hours. To prevent overfitting, the initial learning rates of the 3DVG model and caption head are set to 2e-6 and 2e-4 for ScanRefer [5] (1e-4 and 1e-6 for Nr3D [1]), respectively. We apply learning rate decay at epochs 10 and 20 with a rate of 0.1. As for the input type, we use XYZ coordinates and RGB values as the input, and the number of visual tokens $V$ and query embeddings $Q$ are empirically set to 1024 and 256, respectively. The number of decoder layers $L$ is set to 6. The balancing factors are set default as $\alpha_{1}$ = 1.0 / ($L$ + 1), $\alpha_{2}$ = 5.0, $\alpha_{3}$ = 8.0, $\beta_{1}$ = 5.0, $\beta_{2}$ = 1.0, $\beta_{3}$ = 1.0, $\beta_{4}$ = 0.5 and $\beta_{5}$ = 0.5 for the ScanRefer dataset. For the Nr3D dataset, $\beta_{4}$ and $\beta_{5}$ are adjusted to be 1.0 and 1.0, respectively. During SCST training, we tune the caption head with a batch size of 8 and freeze the rest of the modules in single-GPU for 8 hours. We train 400 epochs with a fixed learning rate of 5e-6 and decay learning rate at epochs 100 and 200 with a rate of 0.1.
Table 1: The 3D referring expression comprehension results on ScanRefer in terms of Acc@0.25IoU and Acc@0.5IoU. Note that what is shown in the table is the optimal accuracy of each method as reported in its paper. Our 3DGCTR surpasses existing state-of-the-art works by a significant margin. EDA-PMB refers to the replacement of single-stage EDA’s [34] visual backbone from Pointnet++ [26] to PointMetaBase [21].
@@ 4.2 Implementation Details
ScanRefer [5] 67.64 49.19 32.06 21.26 38.97 26.10
TGNN [16] 68.61 56.80 29.84 23.18 37.37 29.70
3DJCG [3] 78.75 61.30 40.13 30.08 47.62 36.14
D3Net [6] - 70.35 - 30.50 - 37.87
UniT3D [10] 82.75 73.14 36.36 31.05 45.27 39.14
3DSPS [24] 84.12 66.72 40.32 29.82 48.82 36.98
BUTD-DETR [17] 82.88 64.98 44.73 33.97 50.42 38.60
EDA [34] 85.76 68.57 49.13 37.64 54.59 42.26
EDA-PMB (our backbone) 88.22 73.40 51.76 40.82 57.45 45.91
@@
4.3 Quantitative Comparison
In this subsection, we perform a qualitative comparison on ScanRefer [5] and Nr3D [1] datasets.
3DVG. To evaluate the 3DVG performance, we compare 3DGCTR with several existing 3DVG works in the 3D Referring Expression Comprehension (3DREC) task on ScanRefer [5], involving the state-of-the-art DETR-like methods EDA [34] and BUTD-DETR [17], among others. As evidenced in Table. 1, 3DGCTR consistently exhibits superior performance over these prevailing methods in both two-stage and single-stage settings. This evidence strongly suggests that the newly incorporated caption head has indeed augmented the performance of the 3DVG task. Specifically, by co-training the two tasks, 3DGCTR gets a 0.3% Acc@0.5IoU improvement in overall performance compared to the pre-trained 3DVG backbone EDA-PMB we used. Although there is a slight decrease in the “unique" setting, our 3DGCTR performs better in more complex and “multiple" scenarios (81% of the total testing set), suggesting that the dense caption task can lead to stronger semantic alignment for visual grounding task. The commendable results can be largely attributed to the well-constructed design of our unified 3DVG framework. We also conduct experiments on Nr3D [1], setting the state-of-the-art performance as shown in Table. 4.
3DDC. In 3DDC performance, among the evaluated methods, most employ the standard VoteNet [13] detector, except D3Net [6] and 3DJCG [3]. The state-of-the-art method Vote2Cap-DETR [8] is the first attempt to jointly train 3D Detection and 3D Dense Caption tasks in a single-stage manner, which gains state-of-the-art performance. Table. 2 reports comparisons on ScanRefer [5] dataset. Our model prioritizes enhancing caption accuracy and key content extraction, such as spatial positioning and object attributes, potentially affecting sentence fluidity which leads to a slight performance drop in M and R indicators that focus on the fluency. However, our model excels in the C metric is specially designed for visual description tasks, which offers superior referential value. Specifically, our 3DGCTR achieves 66.11% C@0.5 while Vote2Cap-DETR. [8] achieves 61.81% (4.30% C@0.5↑). Additionally, under SCST, our 3DGCTR outperforms the state-of-the-art method C@0.5 (0.18% C@0.5↑). As shown in Table. 4, a large boost on Nr3D [1] dataset suggests that our pipeline has a greater advantage in dealing with scenes with more complex descriptions.
@@ 4.3 Quantitative Comparison
Scan3Cap [9] 56.82 34.18 26.29 55.27 39.08 23.32 21.97 44.78
MORE [19] 62.91 36.25 26.75 56.33 40.94 22.93 21.66 44.42
SpaCap3d [33] 63.30 36.46 26.71 55.71 44.02 25.26 22.33 45.36
3DJCG [3] 64.70 40.17 27.66 59.23 49.48 31.03 24.22 50.80
D3Net [6] - - - - 46.07 30.29 24.35 51.67
UniT3D [10] - - - - 46.69 27.22 21.91 45.98
Vote2Cap-DETR [8] 71.45 39.34 28.25 59.33 61.81 34.46 26.22 54.40
@@
$\mathcal{X}$-Trans2Cap [39] 61.83 35.65 26.61 54.70 43.87 25.05 22.46 45.28
Scan3Cap [9] - - - - 48.38 26.09 22.15 44.74
D3Net [6] - - - - 62.64 35.68 25.72 53.90
Vote2Cap-DETR [8] 84.15 42.51 28.47 59.26 73.77 38.21 26.64 54.71
@@
Table 3: The 3DVG performance on Nr3D datasets by Acc@0.25IoU as the metric.
Method Acc@0.25IoU
ReferIt3D [1] 35.6
TGNN [16] 37.3
3DSPS [24] 51.5
BUTD [17] 49.1
EDA [34] 52.1
3DGCTR (ours) 52.4
Table 4: The 3DDC performance on Nr3D datasets by CIDer@0.5IoU under MLE and SCST training.
Method MLE SCST
SpaCap3d[33] 33.71 -
$\mathcal{X}$-Trans2Cap[39] - 33.62
D3Net[6] 33.85 38.42
3DJCG[3] 38.06 -
Vote2Cap-DERT[8] 43.84 45.53
3DCGTR (ours) 49.89 52.07
@@ 4.4 Ablation Study
Influence of 3DVG model. We replace the pre-trained 3DVG model with EDA [34] in both single- and two-stage settings. As illustrated in Table. 6, under the joint training setting, various 3DVG backbones achieve enhanced performance in 3DVG tasks (Acc@0.25IoU). Meanwhile, when using the state-of-the-art 3DVG model [34] as our component, we obtain the state-of-the-art 3DDC performance compared to Vote2Cap-DETR [8] (3.29% C@0.5↑), which demonstrate that our “3DDC builds upon 3DVG model" rethinking pipeline offers greater advantages.
Joint Training Scheme. We explore the impacts of different joint training schemes on task performance. After pre-training the 3DVG model without the caption head, if only dense caption data is used for training, the training scheme can be divided into the following three types: 1) only 3DDC: training all components with the same learning rate, 2) only 3DDC (two lr): training all components with different learning rates ($42e-$ 4 and $62e-$ 6), 3) only 3DDC (frozen vg): freezing 3DVG components and training 3DDC components. If both data to jointly train 3DDC and 3DVG tasks, the training scheme can be divided into the following two types: 4) joint training: jointly training 3DDC and 3DVG with the same learning rate, 5) joint training (two lr): jointly training 3DDC and 3DVG with different learning rates ($42e-$ 4 and $62e-$ 6), As displayed in Table. 6, The initial three schemes enhance 3DDC but not 3DVG. The first scheme’s uniform learning rate degrades 3DVG by biasing the model towards DC data, affecting batch normalization statistics. The second and third schemes prevent this by freezing or reducing the 3DVG learning rate, thus maintaining 3DVG accuracy and boosting 3DDC. Conversely, in joint training setting, the fourth scheme’s equal learning rates for both tasks lead to 3DVG overfitting and compromised 3DDC. The fifth scheme’s differential learning rates yield the best 3DDC and improved 3DVG outcomes. Joint training under the fifth scheme shows a 0.3% 3DVG and 1.27% 3DDC performance increase, indicating that a strategic training regimen can effectively co-train 3DDC and 3DVG, mutually enhancing both tasks.
Table 5: Performance on Scanrefer affected by different 3DVG backbone during MLE training. The expressions EDAs and EDAt denote the application of EDA [34] in single- and two-stage settings, respectively. The acronyms 3DGCTR-EDAs and 3DGCTR-EDAt are used to indicate the utilization of EDAs and EDAt as the 3DVG backbone.
@@
Vote2Cap-DETR [8] 61.81 -
EDAs [34] - 53.83
@@
Table 6: Performance on Scanrefer affected by different training schemes during MLE training. ’two lr’ means jointly training 3DDC and 3DVG components with different learning rates, and ’frozen vg’ means only training 3DDC components while freezing the 3DVG backbone.
@@
5) joint training (two lr) 66.11 57.75